# Non-Convex Stochastic Composite Optimization with Polyak Momentum

### v1

@@ 6 Variance reduction effect of momentum
In Section 4, we demonstrated that the vanilla stochastic proximal gradient method cannot converge because the gradient noise variance keeps the iterates away from the stationary point of $F$. In this section, we show that the momentum term in Algorithm 1 can reduce the variance of the gradient noise at the same rate as the gradient norm. This is the key to the convergence of Algorithm 1 without batches. This variance reduction effect has been known in practice and implicitly used in the analysis of [7] and [12]. We precisely characterize such an effect in the composite optimization setting. Proofs are deferred to Appendix B.
We begin by refining the result of Lemma 5.6:
Lemma 6.1.
Given 3.1, 3.3, 3.4, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\frac{3\sqrt{2}L}{(M-L)},a\coloneqq\frac{\sqrt{2}}{8L}$, and $M>(1+3\sqrt{2})L$, we have:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2M}\Delta_{k}+\frac{27\sqrt{2}L}{4M^{2}}\sigma^{2}\,.$ (8)
With Lemma 6.1, we can now precisely quantify the variance reduction effect of the momentum method.
Theorem 6.2.
Given 3.1, 3.3, 3.4, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\frac{3\sqrt{2}L}{(M-L)}$, there exists $M=(1+3\sqrt{2})L+\frac{3\sqrt{3\sqrt{2}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$ such that after at most
$K=\mathcal{O}\left(\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\right)$
iterations of Algorithm 1 it holds that ${\mathbb{E}}\left[\left\lVert\mathbf{m}_{\kappa}-\nabla f(\mathbf{x}_{\kappa})\right\rVert^{2}\right]\leq\varepsilon$, where $\kappa$ is chosen from $\kappa\in\left\{0,\cdots,K-1\right\}$ uniformly at random.
@@ 7 Inexact proximal step
While many existing works (e.g. [16, 14, 39, 17, 37, 42]) rely on the assumption that the proximal step can be solved exactly, this might not always be practically possible [3]. In this section, we briefly discuss extending our analysis to the case where the proximal step is solved inexactly and give an inexactness criterion similar to that of [41].
@@
${\mathbb{E}}\left[\left\lVert\nabla\Omega_{k}(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{M^{2}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\right]}{16}+S_{k}\,,$ (9)
@@
Now we state the convergence result (proofs deferred to Appendix C):
Theorem 7.1.
Given 3.1, 3.3, and 3.4, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\sqrt{\frac{152}{17}}\frac{L}{M-L}$, and at iteration $k$,
${\mathbb{E}}\left[\left\lVert\nabla\Omega_{k}(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{M^{2}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\right]}{16}+S_{k}\,,$
then there exists $M=4L+\sqrt{\frac{8KL\sigma^{2}}{\Phi_{0}}}$ such that after at most
$K=\mathcal{O}\left(\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\right)$
iterations of Algorithm 1 it holds that ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{\kappa})\right\rVert^{2}\right]\leq\frac{\varepsilon}{2}+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}$, where $\mathbf{x}_{\kappa}$ is chosen from $\kappa\in\left\{1,\cdots,K\right\}$ uniformly at random.
In particular, if for any $k$, $S_{k}\leq\frac{\varepsilon}{16}$, then ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{\kappa})\right\rVert^{2}\right]\leq\varepsilon$.
It remains to discuss how to minimize $\Omega_{k}$ such that the inexactness criteria (9) is satisfied. There is a line of research on minimizing gradient norm with SGD variants, e.g. [1], and one can also exploit the structure information in $\psi$ if there is any. Here, we state that SGD suffices for our purpose and give the following proposition, which is a simple modification of Proposition 2.6 in [41]:
Proposition 7.2.
If $\psi$ is convex and $\hat{L}$-smooth, and we have access to an unbiased gradient oracle of $\psi$ with variance at most $\hat{\sigma}^{2}$, then after at most
$T=\mathcal{O}\left((1+\frac{\hat{L}}{M})\ln{\frac{(\hat{L}+M)^{2}}{M^{2}}}+\frac{(\hat{L}+M)\hat{\sigma}^{2}}{MS_{k}}\right)$
@@
The proof of Proposition 7.2 is simple, and we refer interested readers to [41] for the discussions therein.
@@ 8.1 Synthetic quadratic problem
Refer to caption Refer to caption Refer to caption
Figure 1: Comparison of Algorithm 1 and the vanilla stochastic proximal gradient method on the synthetic quadratic problem. The left, middle, and right figures correspond to $\sigma=5,25,125$, respectively. The vanilla stochastic proximal gradient method uses batch sizes $1,5,10$. The x-axis represents the number of gradient samples and is truncated to only show the first $10^{5}$ gradient samples.
@@
We first consider a synthetic quadratic problem inspired by our lower bound construction in Proposition 4.1. We consider the following problem in $\mathbb{R}^{d}$: we set $f(\mathbf{x})=\frac{L\left\lVert\mathbf{x}\right\rVert^{2}}{2}$ and $\psi(\mathbf{x})=\frac{a\left\lVert\mathbf{x}\right\rVert^{2}}{2}$, where $a$ is sufficiently large. Note that the vanilla stochastic proximal gradient method and Algorithm 1 are oblivious to the parameter $a$. We add Gaussian noise to the gradients to control the stochastic level $\sigma^{2}$ of the stochastic oracle. We simulate the batch sample by dividing the variance by the batch size.
@@
In Figure 1, we compare the performance of Algorithm 1 and the vanilla stochastic proximal gradient method. We set $d=5,L=1$ and $a=10^{4}$. We run the vanilla method with batch sizes $1,5,10$. We set $\sigma=5,25,125$ respectively. The parameter $M$ is tuned by a grid search in $\{1,10,100,10^{3},10^{4}\}$ for all methods, and the momentum parameter $\gamma$ is tuned by a grid search in $\{0.1,0.001,10^{-3},10^{-4},10^{-5}\}$. We set the maximum number of iterations to be $10^{4}$, and the tolerance is $0.02$. We see that as $\sigma^{2}$ increases, Algorithm 1 still reaches the desired tolerance, while the vanilla method with batch size $1$ fails to converge in all cases, and with batch sizes $5$ and $10$ only converges when $\sigma=5$. In particular, with batch size $1$, the error of vanilla method oscillates around $36$ with $\sigma^{2}=5^{2}$, around $590$ with $\sigma^{2}=25^{2}$, and around $9375$ with $\sigma^{2}=125^{2}$. In other words, the error of the vanilla method is indeed proportional to $\sigma^{2}$, as predicted by our lower bound result in Proposition 4.1. We also point out that the momentum method exhibits a jump in the error in the first several iterations, consistent with our analysis that the first step of the momentum method incurs an $\mathcal{O}(\sigma^{2})$ error as well.
8.2 Statistical preconditioning experiment
Now we consider the recent statistical preconditioner (proxy training) technique of [17, 41], where for some objective $\ell$, we consider $f=\ell-\hat{\ell}$ and $\psi=\hat{\ell}$. $\hat{\ell}$ is a “statistical preconditioner” defined on a sub-sample of the training dataset. Here we simulate the setup on Cifar-10 dataset [23]. $\ell$ is defined on the whole $50000$ training images, and $\hat{\ell}$ is defined on a subset of $2560$ training images. We follow the implementation of [41], with one difference: at each iteration $k$, [41] computes the full gradient ^$\nabla\hat{\ell}(\mathbf{x}_{k})$ while we only compute a stochastic gradient of batch size $128$. In the experiment, we use batch size of $512$ for $\ell$, and a batch size of $128$ for the SGD updates on $\Omega_{k}$. We perform a grid-search on the parameters $M$ and $\gamma$. The SGD on $\Omega_{k}$ takes $20$ iterations and a step-size $0.01$, which is tuned in [41]. The experiments demonstrate that SGD is effective and reliable for solving the proximal step with sufficient accuracy (see also the experiments in [41]). We see that the momentum method outperforms the vanilla method, and the convergence of the momentum method seems smoother than the vanilla method.
@@ 9 Conclusion
Appendix A Missing proofs in Section 5
We start by giving the proof of Lemma 5.2: See 5.2
@@ Appendix A Missing proofs in Section 5
${\mathbb{E}}\left[\left\lVert\mathbf{m}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$ $={\mathbb{E}}\left[\left\lVert(1-\gamma)\mathbf{m}_{k}+\gamma\mathbf{g}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$={\mathbb{E}}\left[\left\lVert(1-\gamma)(\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k+1}))+\gamma(\mathbf{g}_{k+1}-\nabla f(\mathbf{x}_{k+1}))\right\rVert^{2}\right]$
$=(1-\gamma)^{2}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+{\mathbb{E}}\left[\gamma^{2}\left\lVert\mathbf{g}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$\leq(1-\gamma)^{2}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})+\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\gamma^{2}\sigma^{2}$
Where the in the second equality we used the fact that $\mathbf{g}_{k+1}$ is unbiased. By Young’s inequality, for any $\alpha>0$, we have:
${\mathbb{E}}\left[\left\lVert\mathbf{m}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$ $\leq(1-\gamma)^{2}(1+\alpha){\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\right]$
$\quad+(1-\gamma)^{2}(1+\alpha^{-1}){\mathbb{E}}\left[\left\lVert\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\gamma^{2}\sigma^{2}$
$\leq(1-\gamma)^{2}(1+\alpha){\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\right]$
$\quad+(1-\gamma)^{2}(1+\alpha^{-1})L^{2}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rVert^{2}\right]+\gamma^{2}\sigma^{2}$
where the last inequality follows from the smoothness of $f$. For $\alpha\coloneqq\frac{\gamma}{1-\gamma}$, we have:
${\mathbb{E}}\left[\left\lVert\mathbf{m}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq(1-\gamma){\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\right]+\frac{L^{2}}{\gamma}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rVert^{2}\right]+\gamma^{2}\sigma^{2}$
Now we give the proof of Lemma 5.6:
See 5.6

### v4

@@ 6 Variance Reduction Effect of Momentum
In Section 4, we demonstrated that the vanilla stochastic proximal gradient method cannot converge because the gradient noise variance keeps the iterates away from the stationary point of $F$. In this section, we show that the momentum term in Algorithm 1 can reduce the variance of the gradient noise at the same rate as the gradient norm. This is the key to the convergence of Algorithm 1 without batches. This variance reduction effect has been known in practice and implicitly used in the analysis of Cutkosky and Mehta (2020) and Fatkhullin et al. (2023). We precisely characterize such an effect in the composite optimization setting. Proofs are deferred to Appendix C.
We begin by refining the result of Lemma 5:
Lemma 8.
Let Assumptions 1, 3, 4 and 5 hold, and let $a\coloneqq\frac{\sqrt{2}}{8L}$, $\gamma_{k}\coloneqq\frac{3\sqrt{2}L}{M_{k}-L}$ and $M_{k}>(1+3\sqrt{2})L$ for any $k\geq 0$. Then, for any $k\geq 0$,
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M_{k}}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2M_{k}}\Delta_{k}+\frac{27\sqrt{2}L}{4M_{k}^{2}}\sigma^{2}.$
With Lemma 8, we can now precisely quantify the variance reduction effect of the momentum method.
Theorem 2.
Consider algorithm 1, as applied to solving problem (1) under Assumptions 3, 4 and 5, run for $K=\mathcal{O}\bigl{(}\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\bigr{)}$ iterations with constant coefficients $M_{k}=M=(1+3\sqrt{2})L+\frac{3^{3/2}}{2^{3/4}}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$ and $\gamma_{k}=\frac{3\sqrt{2}L}{M-L}$ for any $10\leq k\leq K-$ 1, where $\Phi_{0}\coloneqq F_{0}+\frac{\sqrt{2}}{8L}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]$, $F_{0}\coloneqq F(\mathbf{x}_{0})-F^{*}$ and $\varepsilon>0$ is a given target error. Then, for the point $\mathbf{x}_{t}$ chosen uniformly at random from $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$ it holds that ${\mathbb{E}}\left[\left\lVert\mathbf{m}_{t}-\nabla f(\mathbf{x}_{t})\right\rVert^{2}\right]\leq\varepsilon$.
@@ 7 Inexact Proximal Step
While many existing works (e.g. Ghadimi et al. (2016); Ghadimi and Lan (2013a); Wang et al. (2019); Hendrikx et al. (2020); Tran-Dinh et al. (2022); Xu and Xu (2022)) rely on the assumption that the proximal step can be solved exactly, this might not always be practically possible Barré et al. (2023). In this section, we briefly discuss extending our analysis to the case where the proximal step is solved inexactly and give an inexactness criterion similar to that of Woodworth et al. (2023).
@@
${\mathbb{E}}\left[\left\lVert\nabla\Omega_{k}(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{M_{k}^{2}}{16}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\right]+S_{k},$ (10)
@@
Now we state the convergence result (proofs are deferred to Appendix D):
Theorem 3.
Consider Algorithm 1, as applied to solving problem (1) under Assumptions 1, 3 and 4, and the approximate stationarity condition at each iteration $k$: ${\mathbb{E}}\left[\left\lVert\nabla\Omega_{k}(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{M^{2}}{16}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\right]+S_{k},$, run for $K=\mathcal{O}\left(\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\right)$ iterations with constant coefficients $M_{k}=M=4L+\sqrt{\frac{8KL\sigma^{2}}{\Phi_{0}}}$ and $\gamma_{k}=\sqrt{\frac{152}{17}}\frac{L}{M-L}$ for any $10\leq k\leq K-$ 1, where $\Phi+0=F_{0}+\sqrt{\frac{19}{156}}\frac{{\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]}{L},F_{0}\coloneqq F(\mathbf{x}_{0})-F^{\star}$ and $\varepsilon>0$ is a given target error. Then, for the point $\mathbf{x}-t$ chosen uniformly at random from $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$ it holds that ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{t})\right\rVert^{2}\right]\leq\frac{\varepsilon}{2}+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}$. In particular, if for any $10\leq k\leq K-$ 1, $S_{k}\leq\frac{\varepsilon}{16}$, then ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{t})\right\rVert^{2}\right]\leq\varepsilon$.
It remains to discuss how to minimize $\Omega_{k}$ such that the inexactness criteria (10) is satisfied. There is a line of research on minimizing gradient norm with SGD variants, e.g. Allen-Zhu (2018), and one can also exploit the structure information in $\psi$ if there is any. Here, we state that SGD suffices for our purpose and give the following proposition, which is a simple modification of Proposition 2.6 in Woodworth et al. (2023):
Proposition 2.
If $\psi$ is convex and $L_{\psi}$-smooth, and we have access to an unbiased gradient oracle of $\psi$ with variance at most $\sigma_{\psi}^{2}$, then after at most
$T=\mathcal{O}\biggl{(}\frac{L_{\psi}+M}{M}\ln\frac{L_{\psi}+M}{M}+\frac{(L_{\psi}+M)\sigma_{\psi}^{2}}{MS_{k}}\biggr{)}$
@@
The proof of Proposition 2 is simple, and we refer interested readers to Woodworth et al. (2023) for the discussions therein.
Refer to caption Refer to caption Refer to caption
Figure 1: Comparison of Algorithm 1 and the vanilla stochastic proximal gradient method on the synthetic quadratic problem. For the vanilla stohastic proximal methods, we also highlight the smoothed curves on top of the original curves that oscillate much more. The left, middle, and right figures correspond to $\sigma=5,25,125$, respectively. The vanilla stochastic proximal gradient method uses batch sizes $1,16,64$. The x-axis represents the number of gradient samples and is truncated to only show the first $10^{5}$ gradient samples.
@@ 8.1 Synthetic Quadratic Problem
We first consider a synthetic quadratic problem inspired by our lower bound construction in Proposition 1. We consider the following problem in $\mathbb{R}^{d}$: we set $f(\mathbf{x})=\frac{L\left\lVert\mathbf{x}\right\rVert^{2}}{2}$ and $\psi(\mathbf{x})=\frac{a\left\lVert\mathbf{x}\right\rVert^{2}}{2}$, where $a$ is sufficiently large. Note that the vanilla stochastic proximal gradient method and Algorithm 1 are oblivious to the parameter $a$. We add Gaussian noise to the gradients to control the stochastic level $\sigma^{2}$ of the stochastic oracle. We simulate the batch sample by dividing the variance by the batch size.
@@
In Figure 1, we compare the performance of Algorithm 1 and the vanilla stochastic proximal gradient method. We set $d=5,L=1$ and $a=10^{4}$. We run the vanilla method with batch sizes $1,16,64$. We set $\sigma=5,25,125$ respectively. The parameter $M$ is tuned by a grid search in $\{10^{0},10^{1},10^{2},10^{3},10^{4}\}$ for all methods, and the momentum parameter $\gamma$ is tuned by a grid search in $\{10^{-1},10^{-2},10^{-3},10^{-4},10^{-5}\}$. We set the maximum number of iterations to be $10^{4}$, and the tolerance is $0.02$. We see that as $\sigma^{2}$ increases, Algorithm 1 still reaches the desired tolerance, while the vanilla method with batch size $1$ fails to converge in all cases, and with batch sizes $16$ and $64$ only converges when $\sigma=5$. In particular, with batch size $1$, the error of vanilla method oscillates around $22$ with $\sigma^{2}=5^{2}$, around $672$ with $\sigma^{2}=25^{2}$, and around $14381$ with $\sigma^{2}=125^{2}$. In other words, the error of the vanilla method is indeed proportional to $\sigma^{2}$, as predicted by our lower bound result in Proposition 1. We also have that for Algorithm 1, $M$ is set to be $10,10,100$ and $\gamma$ is set to be $0.01,0.001,0.0001$ for $\sigma=5,25,125$ respectively, which is consistent with our theoretical prediction that $M$ should increase while $\gamma$ should decrease as $\sigma$ increases. We also point out that the momentum method exhibits a jump in the error in the first several iterations, consistent with our analysis that the first step of the momentum method incurs an $\mathcal{O}(\sigma^{2})$ error as well.
8.2 Regularized Machine Learning Experiment
Now we consider the classical application of composite optimization: regularized machine learning Liu et al. (2015). We use the $\ell_{\infty,1}$ regularizer to regularize the weights on each layer. The proximal step with the $\ell_{\infty,1}$ regularizer is implemented in Murray et al. (2019). We evaluate the performances of Algorithm 1 and the vanilla stochastic proximal gradient method on the Cifar-10 dataset Krizhevsky et al. (2014) with the Resnet-18 He et al. (2016). The regularization parameter is set to be $0.1$, which is observed in Murray et al. (2019) to achive a balance between enforcing sparsity in the model and maintaining the model performance. We use a batch size of $256$ and run $300$ epochs. We use the standard step size parameter $M=10$ (corresponding to a learning rate of $0.1$) for the experiment. We apply a multi-step learning rate scheduler at $150$ epoch and $250$ epoch, with a decay factor of $0.1$. For Algorithm 1, momentum parameter $\gamma$ is set to be $0.1$ by a grid search.
Figure 2: Comparison of Algorithm 1 and the vanilla stochastic proximal gradient method for the $\ell_{\infty,1}$ regularized machine learning problem on Cifar-10 dataset, with Resnet-18. The left and right figures correspond to the training loss and test accuracy, respectively.
We summarize the performances of the two methods in Figure 2. We see that Algorithm 1 outperforms the vanilla method in terms of both training loss and test accuracy. In terms of the test accuracy, Algorithm 1 displays a much smoother curve than that of the vanilla method.
In connection to our theoretical observation that the step size parameter $M$ should increase while the momentum parameter $\gamma$ should decrease as the stochastic level $\sigma$ increases, we observe that, grid searches with respect to the train loss lead to the choices $M=100,100,10$ and $\gamma=0.1,0.1,0.1$ for batch sizes $64,128$ and $256$. It appears that the momentum parameter $\gamma$ is less sensitive to the batch size than the step size parameter $M$, and setting $\gamma=0.1$ might be a good choice in practice. We note that while there is certainly a correlation between the batch size and the stochastic level $\sigma^{2}$, we do not have a direct control over $\sigma^{2}$, as compared to the synthetic problem.
@@ 9 Conclusion
Acknowledgments
We acknowledge partial funding from Helmholtz AI (project Opt4Bio).
Appendix A Additional Experiments
In this section we consider the recent statistical preconditioner (proxy training) technique of Hendrikx et al. [2020], Woodworth et al. [2023], where for some objective $\ell$, we consider $f=\ell-\hat{\ell}$ and $\psi=\hat{\ell}$. $\hat{\ell}$ is a “statistical preconditioner” defined on a sub-sample of the training dataset. Here we simulate the setup on Cifar-10 dataset Krizhevsky et al. [2014]. $\ell$ is defined on the whole $50000$ training images, and $\hat{\ell}$ is defined on a subset of $2560$ training images. We follow the implementation of Woodworth et al. [2023], with one difference: at each iteration $k$, Woodworth et al. [2023] computes the full gradient ^$\nabla\hat{\ell}(\mathbf{x}_{k})$ while we only compute a stochastic gradient of batch size $128$. In the experiment, we use batch size of $512$ for $\ell$, and a batch size of $128$ for the SGD updates on $\Omega_{k}$. We perform a grid-search on the parameters $M$ and $\gamma$. The SGD on $\Omega_{k}$ takes $20$ iterations and a step-size $0.01$, which is tuned in Woodworth et al. [2023]. The experiments demonstrate that SGD is effective and reliable for solving the proximal step with sufficient accuracy (see also the experiments in Woodworth et al. [2023]). We see that the momentum method outperforms the vanilla method, and the convergence of the momentum method seems smoother than the vanilla method.
Refer to caption Refer to caption
Appendix B Missing Proofs in Section 5
We start by giving the proof of Lemma 1: See 1
@@ Appendix B Missing Proofs in Section 5
Indeed,
$\Delta_{k+1}$ $={\mathbb{E}}\left[\left\lVert\mathbf{m}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]={\mathbb{E}}\left[\left\lVert(1-\gamma_{k})(\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k+1}))+\gamma_{k}(\mathbf{g}_{k+1}-\nabla f(\mathbf{x}_{k+1}))\right\rVert^{2}\right]$
$=(1-\gamma_{k})^{2}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+{\mathbb{E}}\left[\gamma_{k}^{2}\left\lVert\mathbf{g}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$\leq(1-\gamma_{k})^{2}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})+\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\gamma_{k}^{2}\sigma^{2}.$
For the second identity, we have used the fact that $\mathbf{g}_{k+1}$ is unbiased. By Young’s inequality, for any $\alpha>0$, we have:
$\Delta_{k+1}$ $\leq(1-\gamma_{k})^{2}(1+\alpha)\Delta_{k}+(1-\gamma_{k})^{2}(1+\alpha^{-1}){\mathbb{E}}\left[\left\lVert\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\gamma_{k}^{2}\sigma^{2}$
$\leq(1-\gamma_{k})^{2}(1+\alpha)\Delta_{k}+(1-\gamma_{k})^{2}(1+\alpha^{-1})L^{2}R_{k}+\gamma_{k}^{2}\sigma^{2},$
where the last inequality follows from the smoothness of $f$. Choosing now $\alpha\coloneqq\frac{\gamma_{k}}{1-\gamma_{k}}$, we get the claimed inequality. ∎
Now we give the proof of Lemma 5:
See 5